# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v3

@@ 5 Experiment
We conduct comparisons based on the SMFG method by excluding the leader-follower update (denoted as S) and the mean field approximation (denoted as MF) individually:
1. SMFG-S: SMFG without leader-follower update. The government agent uses the independent DDPG (de Witt et al. 2020), while interactions among large-scale households are modeled as a mean field game and trained using the mean-field MARL algorithm (Yang et al. 2018).
2. SMFG-MF: SMFG without mean field approximation. Both the government and household agents are trained using a variant of MADDPG algorithm (Lowe et al. 2017), with the government making decisions first, followed by households.
3. SMFG-S-MF: SMFG without both leader-follower update and mean field approximation. In this case, both the government and household agents use independent DDPG.
Macroeconomic Indicators Analysis
Figure 4 presents the results of the ablation studies on various macroeconomic indicators, evaluated across five seeds. The height of each bar represents the mean of the macroeconomic indicators, while the error bars indicate the variance. Our findings reveal that: (1) Excluding either module leads to a significant performance drop compared to the full SMFG method, with this effect becoming more pronounced as the number of agents $N$ increases. This underscores the critical importance of both the S and MF modules. (2) The SMFG-S variant outperforms SMFG-MF and SMFG-S-MF in terms of higher Per Capita GDP, social welfare, and Years, as well as lower Gini index. This suggests the dimensionality reduction of the MF module is crucial for the SMFG method’s overall effectiveness. (3) The low variance of the SMFG method indicates its stability.
Convergence and Social Welfare
Table 2 presents the performance of different baselines across three game theory-related metrics: leader’s payoff, social welfare, and exploitability. Leader’s payoff reflects the performance of the leader agent (i.e., the government), social welfare corresponds to the total utility of all households, and exploitability measures the distance between current policy and Nash equilibrium. The methods for calculating the metrics and optimal values are provided in Appendix D.
From Table 2, we find: (1) SMFG achieves the highest leader’s payoff, the highest social welfare, and near-zero exploitability, indicating that the SMFG policy is closest to both the equilibrium and the maximal social welfare. (2) The high exploitability exhibited by SMFG-S indicates that the leader-follower update module facilitates the algorithm’s convergence toward equilibrium. (3) SMFG-MF outperforms SMFG-S-MF in terms of the leader’s payoff, suggesting that the leader-follower update module aids in optimizing the leader’s objective. However, it is challenging for SMFG-MF to optimize the followers’ social welfare.
@@
SMFG-S-MF -774 -716 0.652 1.023 83 859
SMFG-S -448 -856 3.161 1.213 82 782
SMFG-MF -535 -441 0.782 0.725 54 499
SMFG (ours) 3294 3376 0.002 0.023 98 971
Table 2: Ablation studies of the SMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
Effectiveness Analysis of SMFG
Problem Description
The above experiments have shown the effectiveness of the SMFG method in macroeconomic policymaking. Assuming the government is willing to implement macroeconomic policies learned by the SMFG method, we aim to explore the impact of non-adoption by some households. In our experiments, we introduce Non-SMFG Households whose policies are trained using a behavior cloning algorithm based on real data, reflecting the decision-making styles of real-world households. Details of these behavior cloning experiments are detailed in Appendix E. Additionally, we denote governments and households that adopt the SMFG policy as SMFG Government and SMFG Households, respectively. We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators. The results is shown in Figure 5, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis). GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
Results
From Figure 5, we observe several intriguing findings: (1) SMFG Households outperform Non-SMFG Households across microeconomic indicators such as wealth, income, and utility, suggesting that the policy of SMFG households is more attractive and likely to encourage broader adoption; (2) The presence of Non-SMFG Households does not prevent SMFG Households from achieving near-optimal utility, which is capped at 1 in the TaxAI environment; (3) As the proportion of SMFG Households increases, the per capita GDP also rises, indicating that the government should encourage households to adopt the SMFG policy.
@@ 6 Conclusion
Addressing issues in macroeconomic policymaking, this paper innovatively proposes the SMFG method, which includes dynamic SMFG for modeling and SMFRL algorithm for solving. The SMFG method optimizes government’s policy by considering the dynamic responses of large-scale households to policy shifts, ultimately deriving optimal policies for both government and households. Experimental results demonstrate the SMFG method’s superior performance, stability, and convergence over existing approaches. When extended to real-world scenarios, even if some households do not adopt the SMFG policy, their wealth and utility are not as high as those who do, encouraging more households to adopt the SMFG policy. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the fields of AI for economics and AI for social impact.
@@ Appendix B Assumptions and Limitations
The limitations of our SMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors. (2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games. Currently, our approach is empirically demonstrated by showing that followers converge toward their best responses and that the leader achieves higher performance compared to other baselines. (3) We intend to examine dynamic games between a leader and a large, heterogeneous group of followers, including scenarios where followers dynamically alter their strategies, to determine the optimal leader policy. Addressing these limitations will provide further insights applicable to real-world scenarios.
@@ Appendix C Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows (Saez 2001). The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@ Appendix D Game Theory Metrics
where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.
@@
with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3).
@@
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare (Arrow and Debreu 1954; Malul and Bar-El 2009). Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions defined in Section 3 of the $N$ followers:
@@ Appendix E Additional Results
Algorithm N=100 (hours) N=1000 (hours)
SMFG 4 14
SMFG-MF 3 16
SMFG-S 4 9
SMFG-S-MF 2 6
Free Market 0.25 2
Saez Tax 4 23
AI Economist 6.5 N/A
Table 3: The approximate training times for baselines in our experiments.
Experimental Results of Ablation Studies
@@
In this section, we present additional experimental results from ablation studies, including training curves 6 and a test table 4, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents. We find that the SMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning. More specifically, we compared SMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 6 illustrates the training curves of 4 key macroeconomic indicators under these four settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
@@
Figure 6 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, SMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The SMFG method substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, SMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, SMFG consistently obtains the most optimal solutions across all indicators.
Settings Algorithm Per Capita GDP Social Welfare Income Gini Wealth Gini Years
N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
SMFG-MF 1.23e+05 48.17 0.89 0.93 1.00
SMFG-S-MF 1.21e+05 83.09 0.88 0.92 1.00
SMFG-S 8.66e+07 82.02 0.82 0.83 75.75
SMFG 9.59e+12 96.87 0.52 0.51 300.00
N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
SMFG-MF 1.48e+05 499.01 0.89 0.93 1.02
SMFG-S-MF 1.33e+05 874.53 0.89 0.92 1.00
SMFG-S 1.31e+05 834.93 0.88 0.92 1.50
SMFG 1.10e+13 968.94 0.54 0.53 300.00
N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
SMFG-MF 6.38e+12 93.89 0.57 0.58 268.53
SMFG-S-MF 7.41e+12 98.16 0.53 0.55 300.00
SMFG-S 5.44e+12 98.21 0.50 0.52 300.00
SMFG 1.01e+13 96.90 0.51 0.53 299.89
N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
SMFG-MF 6.82e+12 954.88 0.56 0.62 278.50
SMFG-S-MF 2.79e+12 512.19 0.77 0.81 100.68
SMFG-S 1.13e+05 440.00 0.90 0.93 1.00
SMFG 9.68e+12 975.15 0.52 0.51 300.00
Table 4: Performance metrics of six algorithms across different settings on key macroeconomic indicators.
@@
Training Curves for Various Tax Policies
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 7 displays the training curves and Table 1 shows the test results. Both Figure 7 and Table 1 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.

### v4

@@ 5.1 Experimental Setting
We compare our method against static, dynamic, and game-based policies to validate the necessity of the proposed DSMFG framework (see Table 1). The parameters of the baselines are provided in Appendix G.
  • Static Policies:
    • Free Market [4]: No government intervention.
    • U.S. Federal Tax: The actual progressive personal income tax policy implemented by the U.S. federal government in 2022. This serves as a strong static policy baseline.
  • Dynamic Policies:
    • Saez Tax [50]: A rule-based economic method widely recommended for tax reforms in real world (details in Appendix C).
    • AI Economist [60]: An independent-based policy employing independent Proximal Policy Optimization (PPO), which does not consider multi-agent interactions.
  • Game-based Policies:
    • DSMFG (Ours): Incorporates dynamic Stackelberg Mean Field Games.
    • DSMFG w/o S: DSMFG without Stackelberg structure.
    • DSMFG w/o MF: DSMFG without mean field approximation.
    • DSMFG w/o MF & S: DSMFG excluding both Stackelberg structure and mean field approximation.
5.2 Effectiveness of Dynamic Modeling
Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100. The DSMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
Our DSMFG framework, a dynamic game-based policy, outperforms static, rule-based, and independent policies across key economic metrics (Table 1) and responds more rapidly to economic shocks (Figure 3).
Superior Performance of Dynamic Policies
Table 1 ranks policy performance: DSMFG (dynamic game-based) $\succ$ Saez Tax (dynamic rule-based) $\succ$ U.S. Federal Tax (static) $\succ$ others. In the TaxAI environment, which models complex economic interactions, ineffective policies often trigger termination conditions (e.g., extreme inequality, Gini $>0.9$). Free Market and AI Economist policies, lacking government leadership or multi-agent modeling, fail to achieve sustainable outcomes, as evidenced by the Years metric. Among sustainable policies (U.S. Federal Tax, Saez Tax, DSMFG), dynamic policies outperform static ones: DSMFG achieves a Per Capita GDP 19 times higher than U.S. Federal Tax ($9.59\times 10^{12}$ vs. $4.88\times 10^{11}$) and 4 times higher than Saez Tax ($9.59\times 10^{12}$ vs. $2.34\times 10^{12}$). At $N=1000$, U.S. Federal Tax, based on 2022 static data, fails to sustain economic development, while Saez Tax shows declining Per Capita GDP, Social Welfare, and Gini metrics. In contrast, DSMFG maintains higher Per Capita GDP with comparable welfare and Gini, demonstrating superior scalability and performance over static, rule-based, and AI-based policies.
Rapid Response to Economic Shocks
To assess dynamic responsiveness, we simulate a financial crisis in the TaxAI environment, where all households lose 50% of their wealth at step 100 (see Appendix F.3 for details). As Free Market and AI Economist policies are limited to one-year simulations (Table 1), we compare DSMFG against U.S. Federal Tax and Saez Tax. Figure 3 shows DSMFG (dark blue line) recovering fastest across all metrics: Social Welfare remains stable, Average Income and Consumption recover within 5 steps, and Wealth Gini and Average Wealth stabilize within 15 steps. This unmatched resilience highlights DSMFG’s precise modeling of follower decisions and superior adaptability.
Figure 4: The left subfigure illustrates a scenario with heterogeneous households adopting various policies. The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of DSMFG followers ($N=100$).
5.3 Necessity of Stackelberg and Mean-Field Components
To evaluate the importance of the Stackelberg and mean-field components in DSMFG, we compare DSMFG against its ablated variants (DSMFG w/o S, DSMFG w/o MF, DSMFG w/o MF & S) using macroeconomic metrics (Table 1), training dynamics (Figure 5), and game-theoretic indicators (Table 2). These indicators-leader payoff, exploitability, and social welfare-quantify the leader’s policy optimization, convergence to equilibrium, and follower policy quality, respectively (see Appendix D for details).
Economic performance and training stability
Table 1 demonstrates DSMFG’s superior performance across all economic metrics. Removing either the Stackelberg or mean-field component causes substantial performance degradation. For instance, DSMFG w/o S yields a social welfare of $834.93$ for $N=1000$, a $14\%$ drop from DSMFG, while DSMFG w/o MF reduces social welfare to $499.01$, nearly halving DSMFG’s value. Training curves in Figure 5 further confirm these findings: DSMFG converges stably, while its variants exhibit slower convergence and higher variance (shaded regions, based on five seeds), particularly for DSMFG w/o S and DSMFG w/o MF. These results highlight the indispensable role of both components in achieving robust and optimal policy outcomes.
Table 2: Ablation studies of the DSMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
@@ 5.3 Necessity of Stackelberg and Mean-Field Components
DSMFG (ours) 3294 3376 0.002 0.023 98 971
DSMFG w/o S -448 -856 3.161 1.213 82 782
DSMFG w/o MF -535 -441 0.782 0.725 54 499
DSMFG w/o MF & S -774 -716 0.652 1.023 83 859
Game-Theoretic Analysis
Table 2 provides deeper insights into the contributions of each component. DSMFG achieves a leader payoff of $3294$ for $N=100$ and an exploitability of $0.002$, closely approaching the optimal value of 0. (i) Removing the Stackelberg module (DSMFG w/o S) drastically reduces the leader payoff to $-448$ and increases exploitability to $3.161$, indicating the Stackelberg structure’s critical role in optimizing leader policies and ensuring equilibrium convergence. Similarly, (ii) removing the mean-field component (DSMFG w/o MF) lowers social welfare to $54$ for $N=100$, a $45\%$ reduction from DSMFG’s $98$, underscoring its necessity for effective follower policy optimization. The combined ablation (DSMFG w/o MF & S) reduces the framework to standard multi-agent reinforcement learning, performing comparably to DSMFG w/o MF but with increased exploitability ($1.023$ vs. $0.725$ for $N=1000$). This further emphasizes the Stackelberg module’s importance in large-scale settings, where it facilitates convergence to a stable equilibrium.
5.4 Impact of Mean-Field Homogeneity
DSMFG leverages a mean-field approximation to ensure scalability in modeling and learning, but this introduces an inherent homogeneity assumption. In DSMFG, individual heterogeneity is preserved through state representations, allowing the shared policy to generate personalized actions. This section investigates the impact of shared follower policies by examining: (i) the effects of shared policies on training scalability and efficiency; and (ii) the robustness of DSMFG policies when interacting with heterogeneous follower behaviors.
Scalability and efficiency of shared policy
The shared-policy design in DSMFG significantly improves both scalability and training efficiency compared to heterogeneous-policy variants, such as DSMFG w/o MF and DSMFG w/o MF & S. (i) Scalability: As shown in Table 1, heterogeneous-policy variants suffer from severe training instability, achieving sustainability for only $1$ year and producing lower GDP ($1.48\times 10^{5}$ vs. $1.10\times 10^{13}$ for DSMFG). Additionally, inequality worsens with a higher wealth Gini (0.93 vs. 0.53 at $N=1000$). (ii) Efficiency: Table 3 (Appendix E.1) shows that DSMFG reduces training time by 30% to reach equivalent reward levels. These results highlight the advantages of shared-policy training in large-scale environments, enabling efficient convergence and better economic performance.
Robustness of DSMFG policies
We test DSMFG policy robustness by introducing heterogeneous followers (named Non-DSMFG followers) using behavior-cloned policies derived from real-world data (Appendix E.5). The leader and a subset of followers retain DSMFG-trained policies, while the rest adopt Non-DSMFG followers. We vary the proportion of DSMFG followers (0%, 25%, 50%, 75%, 100%) and track both micro-level (wealth, income, utility) and macro-level (GDP) indicators. Figure 4 reports average values (left Y-axis, bars) and total values (right Y-axis, points).
Results show three key findings: (1) DSMFG followers consistently maintain high utility (96–97) across all proportions, indicating robustness against policy heterogeneity. (2) DSMFG followers outperform Non-DSMFG followers in wealth, income, and utility—often by more than ×$\times 2$—demonstrating the superior effectiveness of DSMFG policies. (3) Per capita GDP increases monotonically with the proportion of DSMFG followers, suggesting that widespread adoption of DSMFG policies can yield substantial macroeconomic gains.
@@ 6 Conclusion
We introduce the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which captures the dynamic, asymmetric, and large-scale nature of government–individual interactions in macroeconomic settings. To solve DSMFG, we develop the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which combines Stackelberg game theory with mean-field approximation to enable scalable and efficient policy learning in large populations. This approach provides a principled and scalable solution to policy optimization problems that are otherwise computationally intractable. Our results underscore the value of integrating game-theoretic modeling with data-driven learning for large-scale economic decision-making. Future directions include extending DSMFG to multi-policy settings, modeling richer behavioral heterogeneity, and calibrating with real-world economic data.
@@ Appendix B Assumptions and Limitations
The limitations of our DSMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors. (2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games. Currently, our approach is empirically demonstrated by showing that followers converge toward their best responses and that the leader achieves higher performance compared to other baselines. (3) We intend to examine dynamic games between a leader and a large, heterogeneous group of followers, including scenarios where followers dynamically alter their strategies, to determine the optimal leader policy. Addressing these limitations will provide further insights applicable to real-world scenarios.
@@ Appendix C Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows [50]. The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@ Appendix D Game Theory Metrics
where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.2.
@@
with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3.2).
@@
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 38]. Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions defined in Section 3 of the $N$ followers:
@@ E.1 Compute Resources
Algorithm Training Time (hours) Utility (years) Utility per training time
N=100 N=1000 N=100 N=1000 N=100 N=1000
DSMFG 4 14 300 300 75.00 21.43
DSMFG w/o MF 3.5 16 1 1.02 0.29 0.06
DSMFG w/o S 4 9 75.75 1.5 18.94 0.17
DSMFG w/o MF & S 2 6 1 1 0.50 0.17
Free Market 0.25 2 1 1 4.00 0.50
Saez Tax 4 23 300 100.58 75.00 4.37
AI Economist 6.5 N/A 1 N/A 0.15 N/A
Table 3: The average training times, utility (Years), and utility per training time for baselines in our experiments. The best values are highlighted in bold, and the second-best values are underlined. Utility is measured using the "Years" metric, which represents the number of simulation steps achievable under a given policy. A higher number of simulation steps indicates better policy performance but also corresponds to increased computational complexity.
E.2 Further Experiments on the Necessity of SMFG
@@
In this section, we present additional experimental results for validating the necessity of SMFG, including training curves Figure 6 and Table 4 and 5, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents. We find that the DSMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning. More specifically, we compared DSMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 6 illustrates the training curves of 4 key macroeconomic indicators under these four settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 and 5 displays the test results of the 7 algorithms across 4 indicators, with each column corresponding to an experimental setting.
@@
Figure 6 and Table 4 and 5 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, DSMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The DSMFG method substantially outperforms other algorithms in solving DSMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, DSMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, DSMFG consistently obtains the most optimal solutions across all indicators.
Table 4: Performance of multiple policies on key macroeconomic indicators for $N=100$ households. The best values are highlighted in bold, and the second-best values are underlined.
Category Subcategory Policies Per Capita Social Income Wealth Years $\uparrow$
GDP $\uparrow$ Welfare $\uparrow$ Gini $\downarrow$ Gini $\downarrow$
Static Policy Non-intervention Free Market 1.37e+05 32.97 0.89 0.92 1.10
Real-data US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
Rule-based Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
Dynamic Policy Independent-based AI Economist 1.26e+05 72.81 0.88 0.91 1.00
without BC Game-based Markov Game 1.21e+05 83.09 0.88 0.92 1.00
Stackelberg Game 1.23e+05 48.17 0.89 0.93 1.00
Mean Field Game 8.66e+07 82.02 0.82 0.83 75.75
DSMFG (ours) 9.59e+12 96.87 0.52 0.51 300.00
Dynamic Policy Independent-based AI Economist 2.03e+12 94.50 0.46 0.48 299.85
with BC Game-based Markov Game 7.41e+12 98.16 0.53 0.55 300.00
Stackelberg Game 6.38e+12 93.89 0.57 0.58 268.53
Mean Field Game 5.44e+12 98.21 0.50 0.52 300.00
DSMFG (ours) 1.01e+13 96.90 0.51 0.53 299.89
Table 5: Performance of multiple policies on key macroeconomic indicators for $N=1000$ households. The best values are highlighted in bold, and the second-best values are underlined.
Category Subcategory Policies Per Capita Social Income Wealth Years $\uparrow$
GDP $\uparrow$ Welfare $\uparrow$ Gini $\downarrow$ Gini $\downarrow$
Static Policy Non-intervention Free Market 1.41e+05 334.79 0.90 0.93 1.00
Real-data US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
Rule-based Saez Tax 6.35e+11 498.88 0.68 0.73 100.58
Dynamic Policy Independent-based AI Economist N/A N/A N/A N/A N/A
without BC Game-based Markov Game 1.33e+05 874.53 0.89 0.92 1.00
Stackelberg Game 1.48e+05 499.01 0.89 0.93 1.02
Mean Field Game 1.31e+05 834.93 0.88 0.92 1.50
DSMFG (ours) 1.10e+13 968.94 0.54 0.53 300.00
Dynamic Policy Independent-based AI Economist N/A N/A N/A N/A N/A
with BC Game-based Markov Game 2.79e+12 512.19 0.77 0.81 100.68
Stackelberg Game 6.82e+12 954.88 0.56 0.62 278.50
Mean Field Game 1.13e+05 440.00 0.90 0.93 1.00
DSMFG (ours) 9.68e+12 975.15 0.52 0.51 300.00
@@
E.3 Training Curves for Various Tax Policies
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 7 displays the training curves and Table 4 and 5 shows the test results. Both Figure 7 and Table 4 and 5 indicate that the DSMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, DSMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the DSMFG.